OpenStudy (anonymous):

graph y=sin(4x-5 pi)+1 over one full period

OpenStudy (mathmale):

Compare y=sin(4x-5 pi)+1 to the more general y = a*sin(bx + c) + d. By doing so, you'll be able to identify the values of a, b, c and d. What do these quantities represent? |a| is the amplitude of the sine function's graph. b is the "frequency." c determines a horizontal phase shift (i. e., the whole sine graph is shifted left or right, according to whether c is positive or negative). d determines a vertical offset (i. e., the whole sine graph is shifted up or down by d units, depending upon the sign of d. Try identifying a, b, c and d, please.

OpenStudy (mathmale):

Once you have b and c, you can then determine the period and horizontal phase shift. The period is (2pi)/b, and the horiz. phase shift is (-c/b().

OpenStudy (anonymous):

he/she is offline. maybe hoping to log back on later with a pat answer ready to go lol

OpenStudy (anonymous):

sorry guys got off to eat lol but i solved it already sorry for the trouble Amplitude=1 Period=pi/2 asymptotes= 5pi/4 and 7pi/4 thanks for the help though

OpenStudy (mathmale):

Glad you were able to do this on your own and cared enough to thank those who had tried to help you.

OpenStudy (anonymous):

haha didnt mean to be rude and log off, but i was hungry lol. :P

OpenStudy (mathmale):

But: Hold it! Where do those asymptotes come from? Does the graph of a sine function have asymptotes?

OpenStudy (anonymous):

well if trying to find one full period it does

OpenStudy (anonymous):

basically a cut off point

OpenStudy (anonymous):

but it doesnt end really

OpenStudy (mathmale):

Careful: "asymptote" doesn't apply to sine or cosine functions. Those trig functions that do have asymptotes are the tan, cot, sec and csc ONLY.

OpenStudy (mathmale):

In the problem you've presented, you have both vertical and horizontal shifts; in the latter case we might call this a "phase shift."

OpenStudy (anonymous):

lol i guess i didnt use the right word :P my teacher told us to use the problem bx-c=0 and bx-c=2pi

OpenStudy (mathmale):

So, what do you think is/are the right word(s)?

OpenStudy (anonymous):

to help find one full period

OpenStudy (anonymous):

well idk cause ive been writing aysmptote on all my papers and she didnt count off

OpenStudy (anonymous):

wouold it be the interval?

OpenStudy (mathmale):

The big question here is: How would you use the results of your analysis to graph a sine function that has both a horizontal (phase) shift and a vertical offset?

OpenStudy (mathmale):

Compare y=sin(4x-5 pi)+1 to the more general y = a*sin(bx + c) + d. By doing so, you'll be able to identify the values of a, b, c and d. What do these quantities represent? |a| is the amplitude of the sine function's graph. b is the "frequency." c determines a horizontal phase shift (i. e., the whole sine graph is shifted left or right, according to whether c is positive or negative). d determines a vertical offset (i. e., the whole sine graph is shifted up or down by d units, depending upon the sign of d. Try identifying a, b, c and d, please. an hour agoDelete 62 mathmale Medals 0 Once you have b and c, you can then determine the period and horizontal phase shift. The period is (2pi)/b, and the horiz. phase shift is (-c/b). (I just cut and pasted material I wrote you earlier.) Please calculate the period and the horizontal phase shift.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

so i would shift everything one space to the right?

OpenStudy (anonymous):

wait its a positive d so i shift the graph one point upward